# MotionScript: Natural Language Descriptions for Expressive 3D Human Motions

### v4

@@ II Related Work
Human Motion Generation. 3D human representation is essential for human-related tasks. For instance, EVA3D [22] and Text2Perform [23] proposed an intermediary pose representation to improve motion synthesis. Moreover, generative models for realistic human motion synthesis including motion prediction [24, 25, 26], motion generation [17, 27], co-speech gesture [28, 29], dance [30], and sign-language generation [4].
Various modalities have been explored for conditional motion generation, including action classes [31, 32], motion descriptions [33, 15, 20], music [34, 35], speech [36, 37], scene context [38, 39], and style [40, 41]. We mainly focus on text-conditioned motion synthesis [33, 15, 42, 43, 14, 18, 20, 19, 44, 5] with a particular emphasis on data augmentation.
@@ II-A Human Motion Semantic Representation
Several methods generate semantic labels or captions from skeleton data. Posebits [55] introduced binary captions describing articulation angle or relative position of joints while [56] used ordinal depth relations of joints as a supervisory signal sourced from human annotators. Poselets [57] extracts intermediate but not easy to interpret pose information considering anthropometric constraints. FixMyPose [58] and AIFit [59] introduced captions offering finer granularity to distinguish poses. Earlier, GDL [60] proposed a rule-based Gesture Description Language (GDL) to represent human body skeleton data with synthetic descriptions. Hierarchical Motion Understanding [61] proposed a program-like representation that described motions with high-level parametric primitives i.e. circular, linear, or stationary and further extended by adding general spline primitives [62]. PoseScript [21] recently introduced a rule-based algorithm that converts pose data into natural language descriptions, starting with lower-level representations such as ‘the left elbow is >$90\degree$ (bent) or <$90\degree$ (straight)’ and translating these into sentences using templates.
@@ III-C Motioncode Extraction
This motioncode categorizes movements into two directions regarding ${\mathbf{sgn}}({\mathbf{M}_{S}})$, with categories ‘right-to-left’, ‘left-to-right’, ‘behind-to-front’, ‘front-to-behind’, and ‘above-to-below’, ‘below-to-above’ corresponding to the X, Y, and Z axes, respectively. Since $\left|{\mathbf{M}_{S}}\right|$ only takes values of $-1$ or $+1$, no predefined intensity classes are assigned.
@@ III-E Motioncode Aggregation
Entity-Based Aggregation merges motioncodes with the same spatial attribute ${\mathbf{M}_{S}}$ but involving different joints of a larger entity. For instance, “The left elbow is getting close to the right foot” and “The left hand is getting closer to the right foot” can be combined as “The left arm is getting closer to the right foot.” This allows joints that are closely related to be described as the motion of a larger entity.
Symmetry-Based Aggregation combines motioncodes with identical ${\mathbf{M}_{S}}$ for symmetric joints on opposite sides of the body. For instance, “the left elbow is bending” and “the right elbow is bending” are aggregated to “the elbows are bending.”
@@
Keypoint-Based Aggregation merges motioncodes that share the same joint set but perform distinct actions. This process factor common key points as the focal subject and aggregate motioncodes both inside a bin and across adjacent bins within a specific range. The subject may be referred as “it” or “they” in the caption while explaining several aggregated motioncodes on that joint set. For example, “the right elbow is significantly bending” and “spreading out from the left elbow” are in the same bin, while “slightly extending” is in the subsequent bin. These are aggregated as the “right elbow is bending and spreading out from the other elbow, afterward, it is slightly extending” with “afterward” indicating the chronological order of motions discussed in Sec. III-E.
Interpretation-based Aggregation fuses the motioncodes with the same spatial attribute ${\mathbf{M}_{S}}$ but operating on different joints. For instance, “the left elbow is slightly bending” and “the right knee is slightly bending” combine to “the left elbow and right knee are slightly bending”.
We apply keypoint and interpretation-based aggregation within a range spanning $T_{range}$ bins before and after, including simultaneous motions within the same bin. The time relation between aggregated motions is maintained to preserve the chronological sequence of actions.
Timecode-Based Aggregation demonstrates the temporal relationship between motioncodes or their chronological order. For instance, consider two motioncodes “the left elbow is slightly bending” followed by “the right knee is spreading out from the other one”. Therefore, merging them while considering their temporal relation would be “the left elbow is bending and immediately after, the right knee is spreading out from the other one”.
Since each aggregated motioncode spans multiple time bins, it’s crucial to preserve their chronological order. For instance, consider “the right elbow is bending” and “the right elbow extends” are in the $n$ and $n+5$ bins respectively, and “the right knee is bending” is in the $n+1$ bin. The aggregated description for the right elbow might be “the right elbow is bending and a few seconds later, it extends” which exceeds the bin for the right knee motioncodes. Therefore, we use “A moment before, the right knee is bending” to reflect the chronological order of motioncodes in the description.
@@ III-F1 Subject Selection
where $(x_{T_{s}},y_{T_{s}},z_{T_{s}})$ and $(x_{T_{e}},y_{T_{e}},z_{T_{e}})$ represent the 3D coordinates of joint $J$ at the start time $T_{s}$ and end time $T_{e}$, respectively. If a specific joint, such as the left hand, contributes more than a certain threshold (e.g., 60%) to the motion, it is identified as the subject (e.g., “the left hand moves away from the right hand.”). Otherwise, we use phrases such as “each other” or “one another” for both joints (e.g., “the left hand and right knee are moving away quickly from each other.”).
@@ III-F2 Pose Injection
We inject the initial pose state and/or final state of posecodes that are relevant to the joints and the type of the motioncodes. For instance, if the motioncode is “the left elbow is slightly bending”, we also add the angle posecodes description “the left elbow is completely extended”. However, relevant posecodes selection depends on their eligibility within the posescript process. There are instances where a posecode may not be eligible, and at times, may require a larger set to cover all necessary posecodes due to aggregation. For instance, the posecode for the left elbow might aggregate to “both elbows are completely bent”, which extends beyond the targeted joint. We then use a modified weighted set cover algorithm [64] to find the maximum covering set of associated posecodes with the minimum number of irrelevant posecodes. Finally, we inject the pose description into the motion description by blending it with transitions selected randomly from the Pose-to-Motion or Motion-to-Pose templates. e.g., {comma, period, ‘and’, ‘from this position’, ‘leads to’, …}.
@@ IV Application: Bridging the gap between LLMs and Motions
An exciting application of MotionScript is its ability to interface with LLMs as a translation layer between high-level, out-of-distribution human motion descriptions and our structured MotionScript format. By converting arbitrary motion descriptions—such as “while following a treasure map, you suddenly find the treasure” or “a person pretending to be an eagle”—into detailed MotionScript captions, LLMs enable a powerful system for generating precise 3D motions from diverse inputs. We prompt an LLM with examples of MotionScript captions to produce MotionScript-like descriptions of the intended motion, as shown in Fig. 1, with additional examples available in the demo video and on the project page.
@@ IV-B Experimental Setup
Caption Test Data. As testing data, we curated a set of 44 out-of-distribution captions, which we call plain captions. These captions were derived from miming and improv acting exercises [^1] https://www.forteachersforstudents.com.au/site/wp-content/uploads/pdfs/tmpl-mime-activity-ideas.pdf. Example captions are shown in Table III and the complete set can be found on the project webpage. In addition, we created detailed captions, which are plain captions converted to a more detailed version of the same text, using ChatGPT. We prompted ChatGPT to either produce detailed captions by prompting it with “Describe a person’s body movements while performing the action {X} in detail. Please respond 2-4 sentences.” (Detailed-LLM), or with examples of MotionScript such as “Describe a person’s body movements while performing the action {X} in detail. Please respond 2-4 sentences in the style of following examples: {motionscript-examples} (see exact prompt on project webpage) (Detailed-MS).
@@ IV-D Experiment 2: LLM Text Augmentation
We randomly selected 34 plain prompts from our set of 44 captions along with their detailed captions, allowing us to compare Detailed-MS + T2M(MS) with Detailed-LLM + T2M(LLM) while maintaining consistency with the trained data for each model. For each caption, participants were asked, “How well does Video X match the prompt?” and provided a rating on a Likert scale from 1 (doesn’t match at all) to 7 (matches extremely well). In addition, they selected their preferred motion for the prompt from the two options. The order of clips was shuffled, and participants could adjust their ratings as needed. Thirty participants completed the survey, and the average duration of this experiment was approximately 25 minutes.
The results showed a strong preference for Detailed-MS + T2M(MS) over Detailed-LLM + T2M(LLM). Detailed-MS + T2M(MS) was preferred in 82% of all n=1817 responses. Likert ratings results from this study are shown in Fig. 5. We can observe a general trend with our proposed method achieving a higher Likert score, with a significantly higher score for 7 captions in particular, which we report in Table III. Overall, through human studies, our experiments confirmed that MotionScript improves text-to-motion generation using our proposed setup.
@@ V Conclusion and Future work
In this work, we introduced MotionScript, a novel semantic representation for 3D human motion that links fine-grained spatial and temporal aspects to natural language descriptions. Unlike traditional captioning methods, MotionScript offers an interpretable mapping between 3D motion and text, enabling large language models to generate 3D motions from high-level descriptions, even for out-of-distribution samples. A human study confirms the effectiveness of our approach. Future work will expand MotionScript to include additional motion features, such as fine motor gestures, facial expressions, and gaze shifts, and compare it to other motion augmentation methods. We also aim to explore its potential in enhancing multimodal AI systems that integrate language, vision, and motion.

### v5

@@ II Related Work
Human Motion Generation. 3D human representation is essential for human-related tasks. For instance, EVA3D [22] and Text2Perform [23] proposed intermediary pose representations to improve motion synthesis. Moreover, generative models for realistic human motion synthesis include motion prediction [24, 25, 26], motion generation [17, 27], co-speech gesture [28, 29], dance [30] and sign-language generation[4]. Various modalities have been explored for conditional motion generation, including action classes [31, 32], motion descriptions [33, 15, 20], music [34, 35], speech [36, 37], scene context [38, 39], and style [40, 41]. We mainly focus on text-conditioned motion synthesis [33, 15, 42, 43, 14, 18, 20, 19, 44, 5] with a particular emphasis on data augmentation.
@@ II-A Human Motion Semantic Representation
Several methods generate semantic labels or captions from skeleton data. Posebits [55] introduced binary captions describing articulation angle or relative position of joints, while [56] used ordinal depth relations of joints as a supervisory signal sourced from human annotators. Poselets [57] extracts intermediate but not easy to interpret pose information considering anthropometric constraints. FixMyPose [58] and AIFit [59] introduced captions offering finer granularity to distinguish poses. GDL [60] proposed a rule-based Gesture Description Language (GDL) to represent human body skeleton data with synthetic descriptions. Hierarchical Motion Understanding [61] proposed a program-like representation that described motions with high-level parametric primitives i.e. circular, linear, or stationary extended by adding general spline primitives [62]. PoseScript [21] recently introduced a rule-based algorithm that converts pose data into natural language descriptions, starting with lower-level representations such as ‘the left elbow is $90\degree$ (bent) or $90\degree$ (straight)’ and translating these into sentences using templates.
@@ III-C Motioncode Extraction
This motioncode categorizes movements into two directions regarding ${\mathbf{sgn}}({\mathbf{M}_{S}})$, with categories ‘right-to-left’, ‘left-to-right’, ‘behind-to-front’, ‘front-to-behind’, and ‘above-to-below’, ‘below-to-above’ corresponding to the X, Y, and Z axes, respectively. Since $\left|{\mathbf{M}_{S}}\right|$ only takes values of 1-1 or +1+1, no predefined intensity classes are assigned.
@@ III-E Motioncode Aggregation
Entity-Based Aggregation merges motioncodes with the same spatial attribute $\mathbf{M}_{S}$} but involving different joints of a larger entity. For instance, “The left elbow gets close to the right foot” and “The left hand gets closer to the right foot” can be combined as “The left arm gets closer to the right foot.” This allows joints that are closely related to be described as the motion of a larger entity.
Symmetry-Based Aggregation combines motioncodes with identical $\mathbf{M}_{S}$} for symmetric joints on opposite sides of the body. For instance, “the left elbow bends” and “the right elbow bends” aggregate to “the elbows bend.”
@@
Keypoint-Based Aggregation merges motioncodes that share the same joint set but perform distinct actions. This process factors common key points as the subject and aggregates motioncodes both inside a bin and across adjacent bins within a specific range. The subject may be referred as “it” or “they” in the caption while explaining several aggregated motioncodes on that joint set. For example, “the right elbow bends significantly” and “spreads out from the left elbow” are in the same bin, while “extends slightly” is in the subsequent bin. These are aggregated as the “right elbow bends and spreads out from the other elbow. Afterward, it extends slightly” with “afterward” indicating the chronological order of motions discussed in Sec. III-E.
Interpretation-based Aggregation fuses the motioncodes with the same spatial attribute $\mathbf{M}_{S}$} but operating on different joints. For instance, “the left elbow bends slightly” and “the right knee bends slightly” combine to “the left elbow and right knee bend slightly”. We apply keypoint and interpretation-based aggregation within a range spanning $T_{range}$ bins before and after, including simultaneous motions within the same bin. The time relation between aggregated motions is maintained to preserve the chronological sequence of actions.
Timecode-Based Aggregation supports the temporal relationship between motioncodes or their chronological order. For instance, consider two motioncodes “the left elbow bends slightly” followed by “the right knee spreads out from the other one”. Therefore, merging them while considering their temporal relation would be “the left elbow bends and immediately after, the right knee spreads out from the other one”.
Since each aggregated motioncode spans multiple time bins, it’s crucial to preserve their chronological order. For instance, consider “the right elbow bends” and “the right elbow extends” are in the $n$ and $n+5$ bins respectively, and “the right knee bends” is in the $n+1$ bin. The aggregated description for the right elbow might be “the right elbow bends and a few seconds later, it extends” which exceeds the bin for the right knee motioncodes. Therefore, we use “A moment before, the right knee bends” to reflect the chronological order of motioncodes in the description.
@@ III-F1 Subject Selection
where $(x_{T_{s}},y_{T_{s}},z_{T_{s}})$ and $(x_{T_{e}},y_{T_{e}},z_{T_{e}})$ represent the 3D coordinates of joint $J$ at the start time $T_{s}$ and end time $T_{e}$, respectively. If a specific joint, such as the left hand, contributes more than a certain threshold (e.g., 60%) to the motion, it is identified as the subject (e.g., “the left hand moves away from the right hand.”). Otherwise, we use phrases such as “each other” or “one another” for both joints (e.g., “the left hand and right knee move away quickly from each other.”).
@@ III-F2 Pose Injection
We inject the initial pose state and/or final state of posecodes that are relevant to the joints and the type of the motioncodes. For instance, if the motioncode is “the left elbow bends slightly”, we also add the angle posecodes description “the left elbow extends completely”. However, relevant posecodes selection depends on their eligibility within the posescript process. There are instances where a posecode may not be eligible, and at times, may require a larger set to cover all necessary posecodes due to aggregation. For instance, the posecode for the left elbow might aggregate to “both elbows bend completely”, which extends beyond the targeted joint. We then use a modified weighted set cover algorithm [64] to find the maximum covering set of associated posecodes with the minimum number of irrelevant posecodes. Finally, we inject the pose description into the motion description by blending it with transitions selected randomly from the Pose-to-Motion or Motion-to-Pose templates. e.g., {comma, period, ‘and’, ‘from this position’, ‘leads to’, …}.
@@ IV Application: Bridging the gap between LLMs and Motions
An exciting application of MotionScript is its ability to interface with LLMs as a translation layer between high-level, out-of-distribution human motion descriptions and our structured MotionScript format. By converting arbitrary motion descriptions—such as “while following a treasure map, you suddenly find the treasure” or “a person pretending to be an eagle”—into detailed MotionScript captions, LLMs enable a powerful system for generating precise 3D motions from diverse inputs. We prompt an LLM with examples of MotionScript captions to produce MotionScript-like descriptions of the intended motion, as shown in Fig. 1, with additional examples in the demo video and project webpage.
@@ IV-B Experimental Setup
Caption Test Data. As testing data, we curated a set of 44 out-of-distribution captions, which we call plain captions. These captions were derived from miming and improv acting exercises [^1] https://www.forteachersforstudents.com.au/site/wp-content/uploads/pdfs/tmpl-mime-activity-ideas.pdf. Example captions are shown in Table III and the complete set can be found on the project webpage. In addition, we created detailed captions, which are plain captions converted to a more detailed version of the same text, using ChatGPT. We prompted ChatGPT-3.5-turbo [66] to either produce detailed captions by prompting it with “Describe a person’s body movements while performing the action {X} in detail. Please respond 2-4 sentences.” (Detailed-LLM), or with examples of MotionScript such as “Describe a person’s body movements while performing the action {X} in detail. Please respond 2-4 sentences in the style of following examples: {motionscript-examples} (see exact prompt on project webpage) (Detailed-MS).
@@ IV-D Experiment 2: LLM Text Augmentation
We randomly selected 34 plain prompts from our set of 44 captions along with their detailed captions, allowing us to compare Detailed-MS + T2M(MS) with Detailed-LLM + T2M(LLM) while maintaining consistency with the trained data for each model. For each caption, participants were asked, “How well does Video X match the prompt?” and provided a rating on a Likert scale from 1 (doesn’t match at all) to 7 (matches extremely well). In addition, they selected their preferred motion for the prompt from the two options. The order of clips was shuffled, and participants could adjust their ratings as needed. Thirty participants recruited via Prolific.com (inclusion criteria: 18+ years old, can read and comprehend English) completed the survey, and the average duration of this experiment was approximately 25 minutes.
The results showed a strong preference for Detailed-MS + T2M(MS) over Detailed-LLM + T2M(LLM). Detailed-MS + T2M(MS) was preferred in 82% of all n=1817 responses. Likert ratings results from this study are shown in Fig. 5. We can observe a general trend with our proposed method achieving a higher Likert score, with a significantly higher score for 7 captions in particular, which we report in Table III. One potential cause of failure cases may be the insertion of parasitic motions in T2M(MS), possibly due to hallucinated MotionScript generated by the LLM. Overall, through human studies, our experiments confirmed that MotionScript improves text-to-motion generation using our proposed setup.
@@ V Conclusion and Future work
In this work, we introduced MotionScript, a novel semantic representation for 3D human motion that links fine-grained spatial and temporal aspects to natural language descriptions. Unlike traditional captioning methods, MotionScript offers an interpretable mapping between 3D motion and text, enabling large language models to generate 3D motions from high-level descriptions, even for out-of-distribution samples. A human study confirms the effectiveness of our approach. Future work will expand MotionScript to include additional motion features, such as fine motor gestures, facial expressions, and gaze shifts, and compare it to other motion augmentation methods. To further improve the readability of the generated descriptions, which are structured but sometimes complex, LLM-based revision could be explored to produce more concise and human-like captions. We also aim to explore its potential in enhancing multimodal AI systems that integrate language, vision, and motion.